# Positional Prompt Tuning for Efficient 3D Representation Learning

### v1

@@ Appendix A Additional Experiments
| K | #TP (M) | PB_T50_RS |
| 1 | 0.9M | 87.40 |
| 2 | 1.1M | 89.52 |
| 3 | 1.4M | 88.06 |
Appendix B Visualization
t-SNE Visualization
In Figure 5, we report the t-SNE visualization results of IDPT(Zha et al. 2023), DAPT(Zhou et al. 2024), and our PPT. Here we report the point cloud feature manifold visualization results on the ScanObjectNN PB_T50_RS dataset. It can be observed that our method, with fewer or on par number of parameters, gets more compact clusters and more separate cluster centers, which means our method gets better results than the parameter-efficient full fine-tuning counterparts.
Attention Visualization
We also report the attention score visualization of either position embedding or patch tokens as the input respectively. Taking the 3D coordinates of patch centers as the input of position embedding, the attention scores of the position embedding input are more holistic, as shown in the top row of Figure 6. That means the position embedding features are concerned with more global information. Meanwhile, the patch token embedding, with all neighbor points around the cluster centers as input, their attention gathered together at some points, as the local counterpart of position embedding inputs. Both concerning global and local features, the point Transformer architectures are deal with complex tasks with a simple MLP as the position encoding module and even get better results than adding those well-designed position embedding methods together, as reported in Table 7.
Appendix C Limitations & Future Work
Although we want to underscore the significance of positional encoding within 3D Transformers and have experimentally validated the use of positional encoding as a form of prompt tuning in classic 3D self-supervised pre-training works such as Point-MAE (Pang et al. 2022) and ReCon (Qi et al. 2023a), we have yet to conduct experimental validation on larger datasets, such as Objaverse (Deitke et al. 2023). Our primary future work will be to experiment with positional prompt tuning on a wider range of 3D foundational models, such as OpenShape (Liu et al. 2023) and ReCon++ (Qi et al. 2024), and to expand its application to a broader spectrum of tasks, including zero-shot learning (Xue et al. 2023; Liu et al. 2023; Qi et al. 2023a, 2024), 3D AIGC (Nichol et al. 2022; Qi et al. 2023b), and 3D LLMs (Qi et al. 2024; Dong et al. 2024), among others (Dong et al. 2022; Peng et al. 2024; Zhang et al. 2023b).

### v2

@@ A.1. Training recipes
To ensure a fair comparison, we employed identical experimental settings to the default fine-tuning method for each baseline. This entails freezing the weights of the pre-trained point cloud backbone and solely updating the newly inserted parameters during training. All experiments are conducted on a single GeForce RTX 3090. In table˜7, we show all our training details.
Appendix B Additional Ablation Study
B.1. Ablation on different settings of MLP elements.
Our PPT contains a number of different MLP components. For the significance of these components, we conducted comprehensive ablation experiments to prove the effectiveness of these parts. As shown in table˜8, we carry out experiments on different structures with the same training settings. Specifically, we test the performance on 3 subsets of Scan_Object_NN dataset (PB_T50_RS, OBJ_BG, OBJ_ONLY). Both Point-MAE and ReCon are considered as baselines. In detail, we choose three different structural settings for ablation: removing the adapter MLP, freezing the weight of the Point MLP during training, and freezing the weight of the Position MLP during training. And our results are listed in the last column. It is worth noting that all other model Settings are consistent with those of PPTwhen these changes are made.
We can tell from the results that all these different changes in model parts result in significant performance reduction, which laterally reflects their effectiveness. Respectively, the addition of the adapter MLP brings the most improvement (3.15% avg. of Point-MAE and 2.65% avg. of ReCon). During fine-tuning, either freezing point MLP (0.77% avg. of Point-MAE and 0.83% avg. of ReCon) or position MLP (0.58% avg. of Point-MAE and 0.90% avg. of ReCon) renders a performance reduction.
| Method | SCAN_OBJECT_NN | no adapter MLP | freezing point MLP | freezing pos MLP | PPT |
| PointMAE (Pang et al., 2022) | PB_T50_RS | 85.39 | 87.89 | 87.93 | 89.00 |
| PointMAE (Pang et al., 2022) | OBJ_BG | 89.85 | 92.94 | 93.12 | 93.63 |
| PointMAE (Pang et al., 2022) | OBJ_ONLY | 90.53 | 92.08 | 92.43 | 92.60 |
| ReCon (Qi et al., 2023a) | PB_T50_RS | 86.22 | 89.10 | 89.04 | 89.52 |
| ReCon (Qi et al., 2023a) | OBJ_BG | 92.08 | 93.46 | 94.49 | 95.01 |
| ReCon (Qi et al., 2023a) | OBJ_ONLY | 91.57 | 92.77 | 91.57 | 93.28 |
Appendix C Additional Visualization
C.1. t-SNE Visualization
In fig.˜6, we report the t-SNE visualization results of Point-MAE(Zhang et al., 2022), DAPT(Zhou et al., 2024), and our PPT. Here, we report the point cloud feature manifold visualization results on the ScanObjectNN PB_T50_RS dataset. It can be observed that our method, with fewer or on par number of parameters, gets more compact clusters and more separate cluster centers, which means our method gets better results than the parameter-efficient full fine-tuning counterparts.